A company reports the following:

Net income $615,000
Preferred dividends $35,000
Shares of common stock outstanding 100,000
Market price per share of common stock $45.82

a. Determine the company's earnings per share on common stock. Round your answer to the nearest cent. Use the rounded answer of requirement a for subsequent requirement, if required.
$

b. Determine the company's price-earnings ratio. Round to one decimal place.
Answer #1

a) Earning per share = (net income-Preferred dividend)/Share outstanding = (615000-35000)/100000 = 5.8 per share

b) Price earning ratio = MPS/EPS = 45.82/5.8 = 7.9 times